In 2002, Home Depot's sales amounted to $58,200,000,000. In 2006, its sales were $90,800,000,000. Write Home Depot's 2002 sales and 2006 sales in scientific notation. Find the increase in sales from 2002 to 2006 in scientific notation. Find what percent that increase is of the 2002 sales in scientific notation.
Answer by checkley77(12569)

HERE'S THE RULE FOR SCIENTIFIC NOTATION:
WHATEVER THE NUMBER YOU START WITH MUST BE CHANGESD SO THERE IS ONE INTEGER BEFORE THE DECIMAL POINT. IF YOU MOVE THEORIGINAL DECIMAL POINT TO THE LEFT TO ACCOMPLISH THIS CHANGE THEN THE NUMBER OF PLACES THE DECIMAL IS MOVED BECOMES THE POSITIVE (+) EXPONENT. IF YOU MOVE THE DECIMAL POINT TO THE RIGHT THE NUMBER OF PLACES BECOMES A NEGATIVE(-) EXPONENT.
EXAMPLES:
.0000000567=5.67*10^-8
5,760,000=5.67*10^6
.00010075=1.0075*10^-4
2,445,000,000,000=2.445*10^12
--------------------------------------------------------------
58,200,000,000=5.28*10^10
90,800,000,000=9.08*10^10
(9.08-5.28)^10^10=3.8*10^10 is the increase in sales.
9.08/5.28=1.446 or 14.46%=1.446*10^1%